# Double robust estimation of functional outcomes with data missing at random

### v2

@@ Appendix B Proofs of the tightness lemmas
$\epsilon^{-2}\mathbb{E}\left(\left\|\xi_{nJ}^{\prime}\right\|^{2}\right)=\epsilon^{-2}\mathbb{E}\left(\left\|n^{-1/2}\sum_{i=1}^{n}X_{iJ}^{\prime}\right\|^{2}\right)\leq\epsilon^{-2}\max_{i}\left(\mathbb{E}\left(\left\|X_{iJ}^{\prime}\right\|^{2}\right)\right),$ (B.1)
@@
Let $r=\sqrt{c/\delta}$ and $c=\sum_{j}^{J_{\epsilon\delta}}\max_{i}\left(\mathbb{E}\langle X_{i},e_{j}\rangle^{2}\right)$. Then the proof is complete by Theorem 7.7.4 in Hsing and Eubank (2015). ∎
B.2 Proof of Lemma 2.2
@@ Appendix C Further results from the Monte Carlo study
We include below simulation results on the OR- and DR-estimators mentioned in the main text, including bias, mean estimated variance as well as Monte Carlo variance, and MSE (combining mean squared bias and Monte Carlo variance). We observe low bias when at least one model is correctly specified. Bias and Monte Carlo variability decrease with sample sizes, and estimated variances correspond to the Monte Carlo variants. Furthermore, Figure C.1 shows the average width of the simultaneous and pointwise confidence bands in one specific scenario and at two different sample sizes. As would be expected, the pointwise bands are notably narrower.
Figure C.1: Average SCB (solid) and PCB (dashed) from 1000 simulation replicates at n = 250 (black) and n = 3000 (green). Based on MVN errors for the OR model with no model misspecification.
Figure C.2: Bias for the Monte Carlo simulations using Gaussian error terms for the OR estimator (top and bottom left panels) and DR estimator (middle and right panels); OR and PS models are outcome and propensity score models. By comparison, the bias of the complete case estimate varies between 0.21 and 1.39.
Figure C.3: Mean estimated variances (solid lines) and Monte Carlo (MC) variance (dashed lines) for the Monte Carlo simulations using Gaussian error terms. Results for the OR estimator (top and bottom left panels) and DR estimator (middle and right panels); OR and PS models are outcome and propensity score models. By comparison, the results for the complete case estimate are at similar levels.
Figure C.4: MSE for the Monte Carlo simulations using Gaussian error terms for the OR estimator (top and bottom left panels) and DR estimator (middle and right panels); OR and PS models are outcome and propensity score models. By comparison, the MSE of the complete case estimate varies between 0.06 and 2.12.
Figure C.5: Bias for the Monte Carlo simulations using multivariate t-distributed error terms for the OR estimator (top and bottom left panels) and DR estimator (middle and right panels); OR and PS models are outcome and propensity score models.By comparison, the bias of the complete case estimate varies between 0.22 and 1.39.
Figure C.6: Mean estimated variances (solid lines) and Monte Carlo (MC) variance (dotted lines) for the Monte Carlo simulations using multivariate t-distributed error terms. Results for the OR estimator (top and bottom left panels) and DR estimator (middle and right panels); OR and PS models are outcome and propensity score models. By comparison, the results for the complete case estimate are at similar levels.
Figure C.7: MSE for the Monte Carlo simulations using multivariate t-distributed error terms for the OR estimator (top and bottom left panels) and DR estimator (middle and right panels); OR and PS models are outcome and propensity score models. By comparison, the MSE of the complete case estimate varies between 0.05 and 2.12.

### v3

@@ B.1 Proof of Lemma 1
$\epsilon^{-2}\mathbb{E}\left(\left\|\xi_{nJ}^{\prime}\right\|^{2}\right)=\epsilon^{-2}\mathbb{E}\left(\left\|n^{-1/2}\sum_{i=1}^{n}X_{iJ}^{\prime}\right\|^{2}\right)\leq\epsilon^{-2}\max_{i}\left(\mathbb{E}\left(\left\|X_{iJ}^{\prime}\right\|^{2}\right)\right),$ (17)
@@
Let $r=\sqrt{c/\delta}$ and $c=\sum_{j}^{J_{\epsilon\delta}}\max_{i}\left(\mathbb{E}\langle X_{i},e_{j}\rangle^{2}\right)$. Then the proof is complete by Theorem 7.7.4 in [14]. ∎
B.2 Proof of Lemma 2
@@ Appendix C Further results from the Monte Carlo study
We include below simulation results on the OR- and DR-estimators mentioned in the main text, including bias, mean estimated variance as well as Monte Carlo variance, and MSE (combining mean squared bias and Monte Carlo variance). We observe low bias when at least one model is correctly specified. Bias and Monte Carlo variability decrease with sample sizes, and estimated variances correspond to the Monte Carlo variants. Furthermore, Figure 3 shows the average width of the simultaneous and pointwise confidence bands in one specific scenario and at two different sample sizes. As would be expected, the pointwise bands are notably narrower.
Figure 3: Average SCB (solid) and PCB (dashed) from 1000 simulation replicates at n = 250 (black) and n = 3000 (green). Based on MVN errors for the OR model with no model misspecification.
Figure 4: Bias for the Monte Carlo simulations using Gaussian error terms for the OR estimator (top and bottom left panels) and DR estimator (middle and right panels); OR and PS models are outcome and propensity score models. By comparison, the bias of the complete case estimate varies between 0.21 and 1.39.
Figure 5: Mean estimated variances (solid lines) and Monte Carlo (MC) variance (dashed lines) for the Monte Carlo simulations using Gaussian error terms. Results for the OR estimator (top and bottom left panels) and DR estimator (middle and right panels); OR and PS models are outcome and propensity score models. By comparison, the results for the complete case estimate are at similar levels.
Figure 6: MSE for the Monte Carlo simulations using Gaussian error terms for the OR estimator (top and bottom left panels) and DR estimator (middle and right panels); OR and PS models are outcome and propensity score models. By comparison, the MSE of the complete case estimate varies between 0.06 and 2.12.
Figure 7: Bias for the Monte Carlo simulations using multivariate t-distributed error terms for the OR estimator (top and bottom left panels) and DR estimator (middle and right panels); OR and PS models are outcome and propensity score models.By comparison, the bias of the complete case estimate varies between 0.22 and 1.39.
Figure 8: Mean estimated variances (solid lines) and Monte Carlo (MC) variance (dotted lines) for the Monte Carlo simulations using multivariate t-distributed error terms. Results for the OR estimator (top and bottom left panels) and DR estimator (middle and right panels); OR and PS models are outcome and propensity score models. By comparison, the results for the complete case estimate are at similar levels.
Figure 9: MSE for the Monte Carlo simulations using multivariate t-distributed error terms for the OR estimator (top and bottom left panels) and DR estimator (middle and right panels); OR and PS models are outcome and propensity score models. By comparison, the MSE of the complete case estimate varies between 0.05 and 2.12.